OpenStudy (emilyjones284):

what is the difference in energy of an electron going from the 4th to the 2nd level in helium?

OpenStudy (jfraser):

plug into the rydberg equation:\[\frac{1}{\lambda} = RZ^2*(\frac{1}{n_1^2} - \frac{1}{n_2^2})\]then use that wavelength to find energy using plank's constant:\[E = h*(\frac{c}{\lambda})\]